A cylinder containing 14.50 L of helium gas at a pressure of 170.0 atm is to be used to fill toy balloons to a pressure of 1.000 atm. Each inflated balloon has a volume of 2.500 L. What is the maximum number of balloons that can be inflated? Report your answer to 1 decimal place. (Remember that 14.50 L of helium at 1.000 atm will remain in the exhausted (empty) cylinder)
1
Expert's answer
2020-11-17T10:38:27-0500

we need to determine the volume of the gas in the cylinder when the pressure is decreased to 1.0atm

write down the known and unknown quantities

P1 = 170.0 atm

V1 = 14.50 L

P2 = 1.0 atm

V2 = ?

Use Boyle's law to find the unknown volume

P1V1 = P2V2


V2 = "\frac{P_1V_1}{P_2}"


V2 = "\frac{170.0atm×14.50L}{1.0atm}"


V2 = 170×14.50 L

V2 = 2465 L


Calculate the number of ballons that can be filled up by dividing the total volume of gas at 1.0 atm by the volume of a single ballon

Nballoon = "\frac{V_2}{V_balloon}"



Nballoon = "\frac{2465}{2.5}"

Nballoon = 986 ballons


Nballons = 0.9×103 balloons